Assigned Problem 1 David Rose Inc. forecasts a capital budget of $500,000 next year with forecasted net income of $400,000. The company wants to maintaina target capital structure of 30 % debt and 70% equity. If the company follows the residual dividend policy, how much in dividends, if any, will it pay? Assigned Problem 2 The following data apply to Elizabeth's Electrical Equipment: All inputs are in millions $20,000 $1,000 $6,000 300 Value of operations Short-term investments Debt Number of shares The company plans on distributing $50 million by repurchasing stock. What will the intrinsic stock price per share be immediately after the repurchase?
Problem Set 7 Assigned Problem 1 NPC is considering either to invest in a project for a new product immediately or 1 year later. If NPC invests in the project today, there will be 75% chance of good market acceptance of the product and 25% chance of bad market acceptance of the product. If market reaction to the new product is good, a cash inflow of $500 million will be realized each year for the next 7 years. If market reaction to the new product is bad, a cash inflow of $25 million will be realized each year for the next 7 years. However, if NPC chooses to wait for 1 year to obtain more information about market tastes, the company would know definitely about the market reaction and would then either proceed with the project or not invest in it at all. The initial cost of the project is $1,500 million. Assuming that all cash flows are discounted at 10%, if NPC chooses to wait a year before proceeding and the project will still end 7 years from now, how much will this increase or decrease the project's expected NPV in today's dollars (i.e., at t 0), relative to the NPV if it proceeds today? a. b. Using the same data, estimate the effect of waiting on the project's risk i.e. calculate by how much the one-year delay will reduce the project's coefficient of variation. (Hint: Use the expected NPV.)
Answer #1

Net income =$ 400,000

Capital budget =$ 500,000

Target capital structure debt = 30%

Equity = 70%

Equity portion of budget = 500000*70% = $350,000.00

So, this amount of $350,000 will be retained from net income of $400,000

Balance income = 400000-350000= $50,000.00

So, Balance $50,000 residual Dividend will be paid.

Note : as per policy 1 different question needs to be answered.